OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = - 1/2 and the point (-2, -3).

OpenStudy (nincompoop):

so first use the point-slope form to solve for the equation in a slope y-intercept form \[y-y_1 = m (x-x_1)\] input the values and then solve for y

OpenStudy (nincompoop):

RESPOND!

OpenStudy (anonymous):

You have to give me a second to think. dang basically the equation is Y-( - 3) = -1/2(x- -2)

OpenStudy (nincompoop):

correct now solve for y

OpenStudy (anonymous):

y=-1/2X-2 is my equation in standard form right?

OpenStudy (nincompoop):

not -1/2x it is -1x/2 or -x/2

OpenStudy (nincompoop):

\[ {y -(-3) = -\frac{1}{2} [x - (-2)]} \rightarrow y+3 = -\frac{1}{2}(x+2) \rightarrow \sf \large {\color {blue} {y+3 = -\frac{x}{2}-1}} \] can you solve for y now?